# INQUIRE-Search: A Framework for Interactive Discovery in Large-Scale Biodiversity Databases

### v2

@@ 2.5.2 Methodology
Figure 12: (Top) Query & Verification. Top INQUIRE-Search returns for "white underside of humpback whale fluke" with incorrect or irrelevant images removed through expert review. Images were marked correct when the right side of the fluke was clearly visible and of sufficient quality for pattern matching. (Bottom) Matching individuals to the HappyWhale database and analysis. A deep learning Re-ID model retrieves candidate matches in the HappyWhale database, which we manually verify, and we then map repeated sightings of confirmed individuals. INQUIRE-Search efficiently surfaces identifiable fluke images from large citizen-science repositories, expanding sighting histories and supporting analyses of movement patterns and habitat use.
We design an approach to efficiently extract identifiable candidate humpback images from the iNaturalist repository. Example retrievals and verifications can be seen at the top of Figure˜12.
  1. 1.
    We queried INQUIRE-Search with a search query including the species name and the body parts required for identification: “white underside of humpback whale fluke”.
  2. 2.
    We analyzed the top 200 images for the search.
  3. 3.
    We manually selected images that provided an unobstructed, clearly visible view of the underside of each whale fluke and were of sufficient quality for analysis. This filtering resulted in 153 images. Most of the 47 excluded images either depicted the wrong part of the whale (e.g., the pectoral fin) or were low quality (blurry or partially obscured).
2.5.3 Results and Analysis
For analysis, we first assess whether any iNaturalist retrievals correspond to known individuals in the HappyWhale dataset. To speed up the matching, we crop each iNaturalist image with Grounding DINO [liu2024grounding] and embed both datasets using a multi-species re-identification model [otarashvili2024multispecies]. For each of the 153 INQUIRE-Search images, we retrieve the top three closest HappyWhale candidates by embedding similarity, then manually review these candidates to confirm matches. Although not conducted by experts, humpback flukes have distinct, easily recognizable patterns that allow reliable identification by non-experts.
@@ 2.5.3 Results and Analysis
Of the 153 candidate images retrieved and verified through INQUIRE-Search, 57 were successfully matched to individuals in the public HappyWhale dataset. Despite inspecting only the top 200 ranked search results, these matches identified 34 unique humpback whales, including multiple sightings of a few individuals, as shown in Figure˜12. These additional sightings extend spatiotemporal coverage of known individuals, demonstrating that INQUIRE-Search can uncover valuable, previously untapped data embedded within large community science repositories.
@@
Figure˜12, which visualizes matched iNaturalist and HappyWhale observations, highlights how this linking expands our ability to track individuals across datasets. Since iNaturalist observations include geolocation metadata, cross-referencing matches provides new spatial information for individuals that were previously represented only in curated catalogs. This capability enables researchers to map additional movement paths, identify unusual sightings, and potentially detect shifts in habitat use when opportunistic images align with structured monitoring efforts.
This proof-of-concept also indicates substantial headroom: evaluating more than the top 200 retrieved images would likely reveal additional high-quality fluke photographs, suggesting that INQUIRE-Search could meaningfully augment re-ID datasets at scale. Importantly, INQUIRE-Search succeeded at retrieving correctly framed, identifiable fluke images despite substantial variation in pose, angle, distance, and environmental settings, which may be conditions under which conventional metadata searches would fail.
2.5.4 Findings
@@
While small in scale, this case study demonstrates that images can be surfaced by INQUIRE-Search, then linked to known individuals using re-ID, providing valuable new information for population monitoring. INQUIRE shows strong potential for retrieving high-quality, correctly posed images that capture identifiable features, which enables the integration of valuable data from previously disconnected sources. However, its broader utility is constrained by the identification step, which still depends on a pre-existing labeled dataset and a trained model for matching. Because INQUIRE cannot assign individual identities on its own, its use is more limited when building datasets from scratch or incorporating previously unseen individuals. In addition, retrieval performance reflects the spatial and behavioral biases of community-sourced photographs, meaning the tool will preferentially surface observations from well-sampled areas and charismatic encounters rather than providing an unbiased representation of whale distribution or behavior.
@@ 3 Discussion
Across our five case studies, we show that INQUIRE-Search accelerates ecological discovery by transforming community-sourced imagery into analyzable evidence. The system enables rapid, expert-guided exploration of ecological questions that would otherwise require extensive field effort or months of manual screening. In each application, INQUIRE-Search helped resolve major data gaps by efficiently recovering information at scale, either through supporting the study of spatiotemporal patterns (bird diet, post-fire regrowth, and phenology), uncovering rare or hard-to-detect events (wildlife mortality), or identifying observations of sufficient quality for downstream analysis (whale re-identification). Together, these examples show how INQUIRE-Search can be used to not only accelerate data discovery but also to fill critical temporal, spatial, and observational gaps across ecological domains.
@@ 3.1 Rethinking the Scientific Process in the Age of Image Search
With the interactive, iterative approach to data discovery enabled by INQUIRE-Search, scientists remain the drivers of inquiry, interpretation, and verification. Incorporating such a tool with an ecological research workflow requires rethinking experimental design. The following discussion highlights alternate parameters to "tune" in the scientific process, from identifying data-driven problem settings with the appropriate scope and iteratively crafting robust queries, to capping verification effort and quantifying uncertainty.
@@ 3.1.1 Which questions are well-posed for INQUIRE-Search?
Determining which scientific questions are best-posed to be addressed with INQUIRE-Search becomes essential. The tool is fundamentally limited by the images that exist in the underlying repository. Questions targeting extremely rare events, under-observed species, or obscure behaviors may yield insufficient data for robust analysis. For example, in the avian diet case study, species with few iNaturalist records produced too few relevant images of feeding behaviors to support quantitative comparisons across time. Conversely, when data are overly-abundant, as in the wildlife mortality study, search still improves efficiency but does not eliminate the need for substantial verification. Additionally, a query is only valuable if the target feature can be unambiguously determined from an image. Prompts such as "robin carrying a worm" or "milkweed with seed pods" are verifiable because the relevant features are visibly explicit. In contrast, prompts such as "bird preparing to migrate" or "tree under stress" are too abstract to evaluate consistently from a still image. Well-posed questions therefore target phenomena that are (1) sufficiently represented in the data pool, (2) not so broad or ubiquitous that they produce large volumes of ambiguous or redundant results, and (3) visually verifiable.
@@ 3.1.2 How to refine a question: iterating on prompts
Search-based discovery depends on clear, specific prompts. Precise language reduces ambiguity in the embedding space and improves the relevance of returned images. Specificity also increases the the verifiability of the outputs: "robin pulling a worm from the soil" will yield cleaner results than "bird foraging," which may return heterogeneous behaviors. Effective prompt-crafting often requires iteration, especially when early results are too broad (refine behavior, context, or anatomical cues) or results contain few matches (broaden the phrasing or adjust the filters). This emphasizes the importance of developing appropriate query design and iterative refinement when using AI-assisted systems like INQUIRE-Search, analogous to testing refining sampling strategies in field studies.
@@ 3.1.4 Identifying Uncertainty and Bias with INQUIRE-Search
INQUIRE-Search inherits the limitations present in the pool of data it retrieves from. Community science platforms such as iNaturalist contain strong observer biases: charismatic or unusual events are overrepresented, while cryptic interactions or remote habitats are under-documented [dimson2023observer]. The accessibility of a study site/subject influences how detectable it is—urban and roadside species are recorded far more often than taxa in remote forests, offshore waters, or high-latitude environments [dimson2023observer]. Methods for observer-effort corrections exist to quantify and mitigate these biases [robinson2018correcting, guilbaultexplorers, mondain2024adaptive, padilla2024assessing], but their effectiveness at a global scale remains limited because they require structured sampling, repeated surveys, or explicit effort measurements that community science data often lack.
Model-driven biases add a second layer of uncertainty. Some prompts may be systematically harder for the embedding model to resolve, such as those that involve examples with low contrast, occlusion, rare species, and unfamiliar backgrounds, introducing an additional layer of uncertainty. VLMs also struggle with prompts requiring complex relational or compositional understanding [alhamoud2025visionlanguagemodelsunderstandnegation, thrush2022winogroundprobingvisionlanguage]. Moreover, recent work shows that VLMs are both miscalibrated (where the expressed confidence does not match true accuracy) [guo2017calibration] and biased (the model defaults to memorized training-set associations instead of analyzing visual evidence) [vo2025vision]. Adapting existing statistical tools for bias estimation and correction [narduzzi2014inverse, geifman2017selective, karimi2023quantifying] could provide a foundation for quantifying uncertainty in discovery-driven search.
@@ 3.2 Expanding Beyond iNaturalist
Although our case studies drew on iNaturalist for its scale and ecological breadth, the open-source design of the INQUIRE-Search codebase makes it adaptable to any large image collection. Early qualitative experiments suggest that the system has potential for other ecological image repositories, such as camera trap datasets. We see INQUIRE-Search less as a single-platform tool and more as a general methodology to enable scientists to get more from their data, flexibly and quickly. Additionally, because the system is open-source and resource-efficient, it has the potential to be deployed widely, enabling scientists, NGOs, and local communities to harness image search for ecological monitoring and hypothesis generation on their own databases.
@@ 4 Conclusion
Growing ecological image databases are a barely-tapped scientific resource, compared to the hidden knowledge they hold. We show that our carefully designed and engineered system, INQUIRE-Search, can enable scientists to work iteratively via open-ended natural language search and verification to rapidly and scalably discover diverse secondary ecological signals, from bird diet to plant phenology. These capabilities point toward a new paradigm of experiment design and data collection in ecology, where hypothesis generation, testing, and data curation occur within a unified, interactive search system over already-collected data. INQUIRE-Search reframes community-sourced imagery from incidental records of species occurrences into an accessible and systematic source of ecological evidence. This perspective democratizes access to data-driven discovery while also raising new challenges around reproducibility, rigor, and inference. It calls on ecologists to integrate search tools into study design: deciding what constitutes a good query, how to measure uncertainty, and how to validate findings across scales. When integrated with careful experimental design, INQUIRE-Search complements traditional ecological methods, offering a scalable path to accelerate discovery, expand participation, and deepen our understanding of biodiversity in a rapidly changing world.
5 Appendix
@@ 5.4 Search Interface
INQUIRE-Search provides a web-based interface (Fig. 13) that surfaces the search index and filters to scientists. The interface includes a query field for natural language queries (e.g., "California condor with a green ‘26’ tag on its wing") and controls for taxonomic, temporal, or geospatial filters. The interface displays a grid of thumbnails of the images returned from the user-provided query, ordered by relevance to the query. The images are clickable, allowing the user to mark images for further analysis. An expanded view of a selected image shows a full-resolution image with complete metadata.

### v3

@@ 3.5 C5: Whale Re-Identification
Figure 11: (Top) Query & Verification. Filtering top INQUIRE-Search outputs for "white underside of humpback whale fluke" where the fluke was clearly visible and of sufficient quality for pattern matching. (Bottom) Matching individuals to the HappyWhale database and analysis. A deep learning Re-ID model retrieves candidate matches in the HappyWhale database, which we manually verify. We then map repeated sightings of confirmed individuals, demonstrating how INQUIRE-Search can support analyses of movement patterns and habitat use.
@@
Of the 153 candidate images retrieved and verified through INQUIRE-Search, 57 were successfully matched to individuals in the public HappyWhale dataset. Despite inspecting only the top 200 ranked search results, these matches identified 34 unique humpback whales, including multiple sightings of a few individuals, as shown in Fig 11. These additional sightings extend spatiotemporal coverage of known individuals, demonstrating that INQUIRE-Search can uncover valuable, previously untapped data embedded within large community science repositories.
@@
Fig 11 also visualizes matched iNaturalist and HappyWhale observations, highlighting how this linking expands our ability to track individuals across datasets. Since iNaturalist observations include geolocation metadata, cross-referencing matches provides new spatial information for individuals that were previously represented only in curated catalogs. This capability enables researchers to map additional movement paths, identify unusual sightings, and potentially detect shifts in habitat use when opportunistic images align with structured monitoring efforts.
Findings
@@
While small in scale, this case study demonstrates that images can be surfaced by INQUIRE-Search, then linked to known individuals using re-ID, augmenting existing datasets and providing valuable new information for population monitoring. INQUIRE shows strong potential for retrieving high-quality, correctly posed images that capture identifiable features, which enables the integration of valuable data from previously disconnected sources. However, its broader utility is constrained by the identification step, which still depends on a pre-existing labeled dataset and a trained model for matching. Because INQUIRE cannot assign individual identities on its own, its use is more limited when building datasets from scratch or incorporating previously unseen individuals. In addition, retrieval performance reflects the spatial and behavioral biases of community-sourced photographs, meaning the tool will preferentially surface observations from well-sampled areas and charismatic encounters rather than providing an unbiased representation of whale distribution or behavior.
@@ 4 Discussion
Across our five case studies, INQUIRE-Search enabled rapid, expert-guided discovery by turning community-sourced images into analyzable ecological data. The system consistently recovered information that would have required substantial fieldwork, helping to resolve major temporal, spatial, and observational gaps.
@@ 4.1 Rethinking the Scientific Process in the Age of Image Search
With the interactive, iterative approach to data discovery enabled by INQUIRE-Search, scientists remain the drivers of inquiry, interpretation, and verification. Incorporating such a tool with an ecological research workflow requires rethinking experimental design. The following discussion highlights parameters to "tune" in this process, from identifying suitable questions and crafting precise prompts, to allocating verification effort and accounting for uncertainty.
@@ 4.1.1 Which questions are well-posed for INQUIRE-Search?
INQUIRE-Search is fundamentally limited by the images that exist in the underlying repository. Questions targeting extremely rare events, under-observed species, or obscure behaviors may yield insufficient data for robust analysis. For example, in the avian diet case study, species with few iNaturalist records produced too few relevant images of feeding behaviors to support quantitative comparisons across time. Conversely, when data are overly-abundant, as in the wildlife mortality study, search still improves efficiency but does not eliminate the need for substantial verification. Useful queries must also be unambiguously visually verifiable: prompts tied to explicit, observable features (“milkweed with seed pods”) are more reliable than abstract states (“tree under stress”). Well-posed questions therefore target phenomena that are (1) sufficiently represented in the data pool, (2) not so broad or ubiquitous that they produce large volumes results, and (3) visually verifiable.
@@ 4.1.2 How to refine a question: iterating on prompts
Search-based discovery depends on clear, specific prompts. Precise language reduces ambiguity in the embedding space and improves the relevance of returned images. Specificity also increases the verifiability of the outputs: "robin pulling a worm from the soil" will yield cleaner results than "bird foraging. Iteration is often necessary, where prompts can be refined by adding behavioral, contextual, or anatomical cues, while overly narrow prompts can be relaxed when returns are sparse.
@@ 4.1.4 Identifying Uncertainty and Bias with INQUIRE-Search
INQUIRE-Search inherits the biases of community-science platforms. Observer preferences skew data toward charismatic, accessible, or unusual events, while cryptic interactions and remote habitats remain under-represented [dimson2023observer]. Although methods for observer-effort corrections to quantify and mitigate these biases exist [robinson2018correcting, guilbaultexplorers, mondain2024adaptive, padilla2024assessing], their effectiveness at a global scale remains limited because they require structured sampling, repeated surveys, or explicit effort measurements that community science data often lack.
Model-driven biases add a second layer of uncertainty. VLMs also struggle with prompts requiring complex relational or compositional understanding [alhamoud2025visionlanguagemodelsunderstandnegation, thrush2022winogroundprobingvisionlanguage]. Moreover, recent work shows that VLMs are both miscalibrated (where the expressed confidence does not match true accuracy) [guo2017calibration] and biased (the model defaults to memorized training-set associations instead of analyzing visual evidence) [vo2025vision]. Adapting existing statistical tools for bias estimation and correction [narduzzi2014inverse, geifman2017selective, karimi2023quantifying] could provide a foundation for quantifying uncertainty in discovery-driven search.
@@ 4.2 Expanding Beyond iNaturalist
Although our case studies drew on iNaturalist for its scale and ecological breadth, the open-source design of the INQUIRE-Search codebase makes it adaptable to any large image collection. Early qualitative experiments suggest that the system has potential for other ecological image repositories, such as camera trap datasets. Because the system is open-source and resource-efficient, it has the potential to be deployed widely, enabling scientists, NGOs, and local communities to harness image search for ecological monitoring and hypothesis generation on their own databases.
@@ 5 Conclusion
Growing ecological image databases remain an underutilized scientific resource. We demonstrate that INQUIRE-Search enables iterative, natural language–driven discovery and verification of diverse ecological signals, supporting scalable analysis of phenomena such as bird diet and plant phenology. This approach suggests a new paradigm in ecological research, in which hypothesis generation, testing, and data curation are integrated within an interactive search system over existing data. By transforming community-sourced imagery from incidental observations into systematic ecological evidence, INQUIRE-Search broadens access to data-driven discovery while raising important questions about rigor, reproducibility, and inference. When combined with careful experimental design, it offers a scalable pathway to accelerate ecological insight in a changing world
Supplementary Information
@@ 5.4 Search Interface
INQUIRE-Search provides a web-based interface (Fig 12) that surfaces the search index and filters to scientists. The interface includes a query field for natural language queries (e.g., "California condor with a green ‘26’ tag on its wing") and controls for taxonomic, temporal, or geospatial filters. The interface displays a grid of thumbnails of the images returned from the user-provided query, ordered by relevance to the query. The images are clickable, allowing the user to mark images for further analysis. An expanded view of a selected image shows a full-resolution image with complete metadata.
@@ 5.6 Data Analysis
5.7 Retrieval and Filtering Results
Species Prompt Template Season Inspected Filtered Rate
(Inv/Vert/Seed/Fruit/Nect/Carr/Plant)
American Robin "<species> with <diet type> in its mouth" Summer 500 per type 338 67.6%
Winter 500 per type 256 51.2%
Red-bellied Woodpecker Summer 500 per type 95 19.0%
Winter 500 per type 161 32.2%
American Tree Sparrow Summer 500 per type 1 0.2%
Winter 500 per type 83 16.6%
Ancient Murrelet Summer 500 per type 0 0.0%
Winter 500 per type 2 0.4%
Gray-cheeked Thrush Summer 500 per type 0 0.0%
Winter 500 per type 0 0.0%
Key: Inv=Invertebrate, Vert=Vertebrate, Nect=Nectar, Carr=Carrion. Counts are placeholders (-/…) where detailed breakdowns were not provided in source text. Total inspected per diet/season/species is 500.
Table 2: Retrieval results for seasonal variation in bird diets. The prompt template is applied to all species. Diet types are abbreviated in the header row.
Target Prompt Filters Insp. Filt. Rate
Young coniferous trees "young coniferous trees in burned forest" Geo: High Park Fire (40.57-40.75 N, 105.18-105.54 W); Date: Post-2012 100 78 78.0%
Young deciduous trees "young deciduous trees in burned forest" 112 45 40.2%
Filters sourced from Table 2.
Table 3: Search parameters and results for post-fire forest regrowth.
Location Prompt Filters Insp. Filt. Rate
Boston (Urban) "dead bird" Geo: Bounding box around Boston, MA; No Taxon filter 1000 543 54.3%
Pioneer Valley (Rural) "dead bird" Geo: Bounding box around Pioneer Valley/Amherst, MA; No Taxon filter 360 79 21.9%
Table 4: Search parameters and results for wildlife mortality.
Stage Prompt Filters Insp. Filt. Rate
Emergence "Milkweed germinating or emerging" Taxon: Asclepias syriaca; Geo: S. Quebec (45.03 to 46.54 N, -74.68 to -71.66 W); No temporal filter 200 45 22.5%
Flowering "Milkweed flowering" 200 169 84.5%
Seeding "Milkweed producing seeds…" 200 161 80.5%
Senescence "Milkweed dying or withering…" 200 52 26.0%
Table 5: Search parameters and results for plant phenology stages.
Target Prompt Filters Insp. Filt. Rate
Humpback whale flukes "white underside of humpback whale fluke" Taxon: Humpback Whale (Megaptera novaeangliae); No Geo/Temp filters 200 153 76.5%
Table 6: Search parameters and results for humpback whale re-identification.